OpenStudy (linstead_cpd_freak):

What is the solution to the equation? 1 - z/5 = 9 A. z = 50 B. z = 44 C. z = –2 D. z = – 40

Mehek (mehek14):

first move 1 to the other side 9 - 1 = ?

OpenStudy (linstead_cpd_freak):

8

Mehek (mehek14):

\(-\dfrac{z}{5}=8\) to get rid of -5, multiply 8 * -5 = ?

OpenStudy (linstead_cpd_freak):

8 x -5 = -40

Mehek (mehek14):

so z = -40 :)

OpenStudy (linstead_cpd_freak):

Ok thank you!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

Mehek (mehek14):

You're welcome ^_^